# Debias your Large Multi-Modal Model at Test-Time with Non-Contrastive Visual Attribute Steering

### v1

@@ 3.2.3 Protected Attribute Mentions
We measure the frequency at which LMMs mention an individual’s protected attributes during open-ended generation, which can be undesirable in circumstances in which commenting on a person’s race, age, gender, or body type would be considered offensive or inappropriate. While our evaluation method can be applied to any social attribute that relates to a set of target tokens, our evaluations focus on four protected attributes of perceived race, age, gender, body type. Due to the variety of ways in which such attributes can be referenced in natural language, we employ two alternative approaches for measuring their frequency: bigram counts and LLM-as-a-judge.
@@
LLaVA-1.5 Steered Dataset Bigram 197 483 549 31 70 282 543 537 274
LLaVA-1.5 Steered Gradient Bigram 583 343 784 109 235 421 650 598 374
@@
Llama-3.2 Steered Dataset Bigram 526 130 629 24 76 120 635 329 456
Llama-3.2 Steered Gradient Bigram 474 190 1240 65 118 337 2380 670 1805
LLaVA-1.5 Unsteered GPT-4 771 1103 235 329 242 1095 21 2689 82
LLaVA-1.5 Steered Dataset GPT-4 365 842 46 158 102 949 1 2269 17
LLaVA-1.5 Steered Gradient GPT-4 783 945 257 301 257 1117 23 2672 68
Llama-3.2 Unsteered GPT-4 869 1641 437 259 246 2169 136 3396 202
Llama-3.2 Steered Dataset GPT-4 893 1564 194 121 125 2800 24 3589 69
Llama-3.2 Steered Gradient GPT-4 792 1312 329 198 162 1872 77 3801 163
Table 1: Frequency of protected attribute mentions per 1k generations for our steered models and other baselines when presented with images from different datasets and open-ended prompts. We find that steering with attributes computed from a dataset nearly always reduces the mention of protected attributes w
@@
We additionally use GPT-4o [2] as a judge to annotate the amount of attribute-related text in each generation. Using a two-shot prompt with OpenAI’s Structured Output API, GPT-4o returns both the count of protected attributes and their corresponding spans. This method has proven to be highly reliable, with minimal discrepancies between the reported counts and the identified spans. Our manual inspection of GPT-4o’s highlighted spans confirmed that it captures a broad, but justified, set of terms that refer to the target attributes.
@@ 3.2.4 Sentiment analysis
Another potential form of bias in LMMs is a disparity across groups in the sentiment of generated text sequences. To investigate whether our steering approach can reduce such disparities, we use VADER [20], a popular lexicon and rule-based sentiment analysis tool [37, 42]. Specifically, we report the compound score produced by VADER across different groups of perceived race, gender, age, and body type attributes in our evaluated datasets. A compound score $>=0.05$ indicate positive sentiment while scores $<=-0.05$ indicate negative sentiment.
@@ 4.1 Experimental details
We conduct experiments by generating responses from two LMMs with and without our steering methods: LLaVA 1.5 [23] and Llama-3.2-Vision-Instruct [12]. We generated a total of 1,820,904 responses for the five prompts, three datasets, two LMMs, and three methods evaluated in our experiments. Our generation experiments were conducted on 512 Intel® Gaudi 2 AI accelerators in the Intel® Tiber AI Cloud.
@@
For each LMM, we compute steering directions by computing the normalized difference between contrasting activations from biased and benign text generations. Biased text is generated from a specific prompt applied to 256 image samples, while benign text is sourced from the LLaVA-Instruct-80K dataset [24] by excluding instances with protected attributes. We evaluate 32 candidate ablation directions based on a held-out set of 10 image-prompt pairs, selecting the most effective direction for further experiments.
@@ 4.2 Protected attribute evaluation
Table 1 shows our main results for evaluating the effectiveness of steering LMM generations from mentioning protected attributes. Across all three datasets, we see that on average and across nearly all conditions, steering significantly reduces the LMM’s propensity to generate text related to the target attributes. This result holds for both the LLM-as-judge as well as the bigram count evaluation methods. One clear trend is that as the model capability increases, so does the effectiveness of our Steered Gradient steering approach. While Steered Dataset is nearly always more effective for the 7B LLaVA model, the optimization-based method is more viable for Llama-3.2-11B. What’s more, Steered Gradient is an improvement over Steered Dataset in the few cases where Steered Dataset shows no improvement over the unsteered baseline. As mention frequency of protected attributes is the most straightforward metric on our approach for debiasing, we believe this shows a significant improvement over the baseline, with an overall improvement of 29% for our Steered Dataset method.
@@ 4.3 Group-specific sentiment
LLaVA-1.5 Steered Dataset 41.6 14
LLaVA-1.5 Steered Gradient 16.4 11
@@
Llama-3.2 Steered Dataset 169.2 29
Llama-3.2 Steered Gradient 309.9 49
Table 2: Variance and range of negative sentiment scores per 1k generations, measured across perceived race groups using the SocialCounterfactuals dataset.
To investigate the impact of our steering methods on differences in the sentiment of text produced for different protected attribute groups, we calculated the frequency of negative text detections per 1k generations for each perceived race group annotated in the SocialCounterfactuals dataset. Table 2 provides the variance and range of the resulting values, calculated across race groups. Higher values of variance and range indicate a greater degree of variability in the frequency of negative text generation across different groups. We observe that both of our steering approaches achieve reductions in this measure of bias with LLaVA-1.5, with our Steered Gradient method providing a particularly large reduction in negative sentiment variability. Additionally, our Steered Dataset approach also reduced negative sentiment variability for Llama-3.2. These results point to additional debiasing effects produced by our steering method beyond simply reductions in the frequency of protected attribute mentions.
@@ 4.4 Occupation mentions
LLaVA-1.5 Steered Dataset -19.43% -17.38% -30.12% -13.18% -2.77% -4.50% -18.34%
LLaVA-1.5 Steered Gradient 0.46% -6.23% -0.08% 1.85% -0.69% 0.00% -3.81%
Llama-3.2 Steered Dataset -5.48% -15.02% -7.84% -2.83% 42.37% 38.55% 4.20%
Llama-3.2 Steered Gradient -3.96% -0.40% 2.90% 8.53% -9.86% 7.25% -7.63%
Table 3: Comparison of Steering on open ended accuracy mentions with LLaVA and Llama LMMs relative to their unsteered baselines for SocialCounterfactuals and FACET.
In table 3 we measure how often the steered models correctly mention the occupation of the person in the image in our open ended prompt generation experiments. While we find the occupation mentions often go down, there is a high variance across the experimental settings, with the interesting result for Llama-3.2 on FACET significantly increasing occupation mentions. We also find the Steered Gradient method to stay more similar in its overall generations to the baseline, hence its lower changes with respect to the unsteered models.
4.5 Accuracy of steered models
We employed the LLM-as-a-judge approach [46] to investigate whether steering affects the accuracy of generated responses. We used GPT-4o to evaluate whether LLaVA’s text responses, with and without steering, match the corresponding image. GPT-4o was given the image and prompt: “Does the description match the image? Answer with Yes or No.” Manual analysis showed that GPT-4o responds “No” when the generation contains extra details not present in the image. The results show no significant difference in accuracy between baseline and steered LLaVA models, indicating that steering does not degrade performance.
@@ 5.1 Impact of steering on token probabilities
Figure 3: The global changes in probabilities of predicting given tokens on the age subset of FACET. The table reflects generated outputs from 100 samples, sorted by most changed, showing how overall steering only effects the raw probabilties in a sublte way.
Figure 3 shows the effectiveness of steering techniques in reducing bias in LMM token predictions. After intervening to ablate biased directions in the model’s internal representations, we observe a shift toward more neutral, contextually appropriate tokens, with biased terms related to protected attributes being suppressed. This effect is consistent in both single-image examples and across 100 generations from the FACET age split.
5.2 Loss Function Analysis for One Step Gradient-Based Optimization
@@
Mean Squared Error Loss with Target Logits

### v2

@@ 3.2.2 Protected attribute mentions
As our primary metric, we measure the frequency at which LMMs mention an individual’s protected attributes during open-ended generation, which can be undesirable in circumstances in which commenting on a person’s race, age, gender, or body type would be considered offensive or inappropriate. While our evaluation method can be applied to any social attribute that relates to a set of target tokens, our evaluations focus on four protected attributes of perceived race, age, gender, body type. Due to the variety of ways in which such attributes can be referenced in natural language, we employ two alternative approaches for measuring their frequency: bigram counts and LLM-as-a-judge.
@@
LLaVA-1.5 Reprompting Bigram 957 708 1405 164 415 759 1211 275 392
LLaVA-1.5 Explanation Bigram 1688 1351 2627 318 795 1639 2198 197 253
LLaVA-1.5 PAR Bigram 197 483 549 31 70 282 543 537 274
LLaVA-1.5 PARzero Bigram 583 343 784 109 235 421 650 598 374
@@
Llama-3.2 Reprompting Bigram 301 211 875 47 147 357 2640 856 617
Llama-3.2 Explanation Bigram 231 196 848 48 92 352 957 1663 1351
Llama-3.2 PAR Bigram 526 130 629 24 76 120 635 329 456
Llama-3.2 PARzero Bigram 474 190 1240 65 118 337 2380 670 1805
LLaVA-1.5 Unsteered GPT-4o 771 1103 235 329 242 1095 21 2689 82
LLaVA-1.5 Reprompting GPT-4o 592 933 153 182 207 1475 80 2639 127
LLaVA-1.5 Explanation GPT-4o 546 962 186 173 240 1072 75 2521 162
LLaVA-1.5 PAR GPT-4o 365 842 46 158 102 949 1 2269 17
LLaVA-1.5 PARzero GPT-4o 783 945 257 301 257 1117 23 2672 68
Llama-3.2 Unsteered GPT-4o 869 1641 437 259 246 2169 136 3396 202
Llama-3.2 Reprompting GPT-4o 326 961 126 85 326 961 126 1180 125
Llama-3.2 Explanation GPT-4o 295 694 216 100 295 694 216 488 155
Llama-3.2 PAR GPT-4o 893 1564 194 121 125 2800 24 3589 69
Llama-3.2 PARzero GPT-4o 792 1312 329 198 162 1872 77 3801 163
Table 1: Frequency of protected attribute mentions per 1k generations. We evaluate our PAR/PARzero approaches as well as the baselines when evaluated on images from different datasets and open-ended prompts. Both steering methods improve on the baselines in most cases. PAR reduces the mention of protected attributes the most, at the cost of reduced efficiency compared to PARzero.
@@
We additionally use GPT-4o [2] as a judge to annotate the amount of attribute-related text in each generation. Using a two-shot prompt with OpenAI’s Structured Output API, GPT-4o returns both the count of protected attributes and their corresponding spans. This method has proven to be highly reliable, with minimal discrepancies between the reported counts and the identified spans. Our manual inspection of GPT-4o’s highlighted spans confirmed that it captures a broad, but justified, set of terms that refer to the target attributes. See the supplementary material for details.
@@ 3.2.3 Sentiment analysis
An important effect of bias in LMMs is a disparity across groups in the sentiment of generated text sequences. To investigate whether our steering approach can reduce such disparities, we use VADER [20], a popular lexicon and rule-based sentiment analysis tool [36, 42]. Specifically, we report the compound score produced by VADER across different groups of perceived race, gender, age, and body type attributes in our evaluated datasets. A compound score $\geq 0.05$ indicate positive sentiment while scores $\leq-0.05$ indicate negative sentiment.
@@ 4.1 Experimental details
We conduct experiments by generating responses from two LMMs with and without our steering methods: LLaVA 1.5 [23] and Llama-3.2-Vision-Instruct [11]. We generated a total of 1,820,904 responses across the five open-ended prompts, three image datasets, two LMMs, two steering methods and three baselines evaluated in our experiments. Our generation experiments were conducted on 512 Intel® Gaudi 2 AI accelerators in the Intel® Tiber AI Cloud.
Baseline approaches
We evaluate three baseline methods against our steering approaches. Specifically, we use the two approaches proposed by Gallegos et al. [13]: self-debiasing via explanation (explanation), and self-debiasing via reprompting (reprompting). Both methods work by prompting an LLM to regenerate an output, but to do so while removing any bias. Additionally, the explanation method prompts the LLM to list any biases present in its primary response, and then correct for them in a second attempt. As a third baseline, we simply evaluate unsteered versions of the LLaVA and Llama LMMs.
@@
For PAR, we compute steering directions as described in section (3.1) with respect to contrasting activations from generations containing protected attribute mentions, and benign text generations. To obtain generations that mention protected attributes, we pair targeted prompts with 256 synthetic images. For benign text we sample 256 image-prompt pairs from the LLaVA-Instruct-80K dataset [24], excluding instances with protected attributes. For PARzero we only make use of the input image, as well as a list of attribute-related tokens to construct the steering vector. For both PAR and PARzero we follow prior work [37, 34, 25] by computing the steering direction with respect to a layer near the middle of the language model. Further details of the experimental design can be found in section (8) of the supplementary material.
@@ 4.2 Protected Attribute Evaluation
Table (1) presents our main results for evaluating how well our proposed steering methods reduce a LMM’s tendency to mention protected attributes. Across all three datasets, both PAR and PARzero lower both LLaVA and Llama’s propensity to generate target-attribute text compared to the unsteered baseline, as confirmed by both the LLM-as-judge and bigram count metrics. PAR and PARzero also improve over prompt-based baselines in nearly all conditions under the bigram count metric, as well as in all conditions with LLaVA under the LLM-as-judge metric. We observe that Llama-3.2 tends to respond well to the prompt-based baselines under the LLM-as-judge metric, especially for synthetic data e.g. SocialCounterfactualsWe hypothesize that the extensive chat tuning of Llama (relative to Llava 1.5) enables it to respond in a manner that the satisfies the GPT4o judge. We also believe that steering Llama-style LMMs may present additional challenges related to their per-layer vision-language fusion. Finally, we note that PARzero improves over the “Unsteered” baseline in nearly all conditions. PARzero does not require iterative prompting like the baselines, or an offline computation stage like PAR, making it the most efficient option for reducing protected attribute mentions in LMMs. We find that PARzero makes the greatest gains over the “Unsteered” baseline real data (FACET, DA-COCO) with Llama-3.2. This possibly signals that stronger models like Llama-3.2 PARzero is a
@@ 4.3 Group-specific Sentiment
LLaVA-1.5 Reprompting 70.8 64
LLaVA-1.5 Explanation 25.7 19
LLaVA-1.5 PAR 41.6 14
LLaVA-1.5 PARzero 16.4 11
@@
Llama-3.2 Reprompting 52.53 48
Llama-3.2 Explanation 188.1 152
Llama-3.2 PAR 169.2 29
Llama-3.2 PARzero 309.9 49
Table 2: Variance and range of negative sentiment scores. We compute each score per 1k generations, measured across perceived race groups using the SocialCounterfactuals dataset.
To investigate the impact of our steering methods on differences in the sentiment of text produced for different protected attribute groups, we calculated the frequency of negative text detections per 1k generations for each perceived race group annotated in the SocialCounterfactuals dataset. table (2) provides the variance and range of the resulting values, calculated across race groups. Higher values of variance and range indicate a greater degree of variability in the frequency of negative text generation across different groups. We observe that both of our steering approaches achieve reductions in this measure of bias with LLaVA-1.5, with PARzero providing a particularly large reduction in negative sentiment variability. Additionally, PAR also reduced negative sentiment variability for Llama-3.2. These results point to additional debiasing effects produced by our steering method beyond simply reductions in the frequency of protected attribute mentions.
@@ 4.4 Occupation Mentions
LLaVA-1.5 PAR -19.43% -17.38% -30.12% -13.18% -2.77% -4.50% -18.34%
LLaVA-1.5 PARzero 0.46% -6.23% -0.08% 1.85% -0.69% 0.00% -3.81%
Llama-3.2 PAR -5.48% -15.02% -7.84% -2.83% 42.37% 38.55% 4.20%
Llama-3.2 PARzero -3.96% -0.40% 2.90% 8.53% -9.86% 7.25% -7.63%
Table 3: Comparison of PAR and PARzero on open-ended occupancy mentions relative to their unsteered baselines. We compare each method on SocialCounterfactuals and FACET. Large relative changes in occupancy mentions means that we are altering the information content of the response, which may not be desirable. PARzero tends to retain the occupation content present in the baseline response better than PAR.
In table (3) we show how often the steered models correctly mention the occupation of the person in the image in our open-ended prompt generation experiments, relative to the unsteered baseline. A higher absolute difference from the baseline indicates that the information content of the response has changed. Depending on the application it may be desirable for the generation to stay close to the unsteered baseline while removing protected attribute mentions. In this regard, PARzero outperforms PAR by having little change in occupancy accuracy relative to the baseline. For PAR, while we find the occupation mentions often decrease, there is a high degree of variability across the experimental settings; interestingly, for PAR with Llama-3.2, correct occupation mentions on FACET significantly increase relative to the baseline.
4.5 Fidelity of Steered Models to Image Content
Given the previous results, its an open question if general captioning capabilities are degraded by steering. To investigate this we employed the LLM-as-a-judge approach [46] to evaluate whether steering affects the accuracy of generated responses. We used GPT-4o to evaluate whether LLaVA’s text responses, with and without steering, match the corresponding image. GPT-4o was given the image and prompt: “Does the description match the image? Answer with Yes or No.” Manual analysis showed that GPT-4o responds “No” when the LMM hallucinates image details. The results in table (4) show no significant difference in accuracy between baseline and steered LLaVA models, indicating that steering does not degrade performance.
Method SocialCounterfactuals DA-COCO
Unsteered $70.53\%$ $64.47\%$
PAR $71.77\%$ $64.33\%$
Table 4: Steering captioning performance. Percentage of LLaVA generations (%) evaluated by GPT-4o as matching the corresponding image for images depicting perceived race.
@@ 5.1 Impact of steering on token probabilities
Figure 3: Global changes in probabilities of predicting given tokens on the age subset of FACET. The table reflects generated outputs from 100 samples, sorted by most changed, showing how overall steering only effects the raw probabilties in a sublte way.
Figure (3) shows how steering techniques can subtly augment the token sampling distribution of the LMM to achieve the desired behavior modification, in this case, reducing the frequency of protected attribute mentions. After intervening to remove biased directions in the model’s internal representations, we observe a shift toward more neutral language, with tokens related to protected attributes being suppressed. We observe this effect in single-image examples as well as in aggregate. We show non-cherry-picked individual examples of steering behavior in section (14) of the supplementary material.
5.2 Impact of PARzero Adaptive Step-size
In section (3.1) we introduced an adaptive method for optimizing the visual representation toward a target attribute. We chose an adaptive target for our loss calculation on the intuition that setting a fixed value would not generalize to multiple attributes. The PARzero loss should result in increased sampling probabilities for tokens related to the target protected attribute. Depending on the training of the LMM, the chosen protected attribute, and the input image, the optimization target i.e. how much we need to increase token probabilities could differ greatly. To show this, we perform an ablation study over the optimization target ($\epsilon$), where we compare fixed $\epsilon$ values to our adaptive method. Figure (4) shows that our adaptive approach performs better than using a fixed value by a wide margin.
Figure 4: Comparing adaptive vs. fixed loss targets for PARzero. We evaluate on 100 samples of the perceived race subset of the FACET dataset.
6 Discussion
Our findings underscore the substantial benefits of an efficient, training-free technique for model steering in large multi-modal models. We showed that PAR and PARzero can effectively guide the model to remove specific protected attribute text, such as perceived age, race, or gender, from its outputs. Given the rapid deployment of LMMs across high-impact sectors, our approach represents a scalable and impactful method for aligning models with ethical and societal expectations without extensive retraining or modification. The elimination of a need for full model re-training or fine-tuning directly translates to significant savings in both energy and hardware resources. As AI practitioners grapple with the environmental impact of increasingly large models, our method contributes a sustainable alternative. This not only allows for quick model refinement but also aligns with the broader goals of sustainable AI development, where resource-intensive operations can be minimized without sacrificing efficacy.
Limitations
Despite our best efforts to improve the fairness of generative AI models, we acknowledge that our choice of models, methodologies, and datasets may themselves contain latent biases which limit our ability to address this multi-faceted problem. The datasets employed in our evaluations as well as our automated evaluation methods which employ LLMs may themselves contain biases which could influence our results. Nevertheless, we feel that our manual validations and the scale of the study we conducted necessitated the use of these resources. We also acknowledge that our study was only conducted with the English language. Additional work is needed to explore the effectiveness of our steering methods for other languages. Despite these limitations, we hope that the findings of our study will raise awareness of the need for methods to debias LMMs and promote the development of more responsible and fair generative AI models.
Future work
Finally, our single-gradient-step approach PARzero offers a promising framework for further development in bias mitigation and model alignment. This initial study suggests a strong foundation for rapid intervention in model behavior, which could be extended to other forms of model control. We also see an opportunity to explore techniques for steering different LMM architectures, given the different performance of LLaVA vs Llama. With the success observed in attribute suppression, further research may explore other types of steering or alignment tasks that could similarly benefit from a minimally invasive, gradient-based intervention. Next steps should aim to expand bias mitigation techniques to encompass a broader spectrum of attributes and assess the long-term impacts of steering interventions on model performance.
Supplementary Material
7 PARzero Loss Analysis
@@
Mean squared error with target logits